# Gravothermalizing into primordial black holes, boson stars, and cannibal stars

### v2

@@ Abstract
Very little is known about the cosmological history from after the end of inflation until Big Bang Nucleosynthesis. Various well-motivated models predict that the universe could have undergone a period of matter domination in this early epoch. We demonstrate that if the particles causing matter domination have self-interactions, they can form halos that undergo a gravothermal collapse. We thus propose a novel scenario for the formation of primordial black holes, which in particular can lie within the asteroid-mass range. We also find that it is not only black holes that can form in the aftermath of a gravothermal evolution. We show that number-changing annihilations of the particles can create sufficient heat to halt the gravothermal evolution, thus forming a “cannibal star”. Likewise, the pressure from the particle’s repulsive self-interactions can form a boson star during a gravothermal evolution. Thus, our study highlights that structure formation in the early universe can have a rich phenomenology.
@@ I Introduction
The period of matter domination in the standard Hot Big Bang history lasts about 8 $e$-folds. An EMDE, on the other hand, can have a much longer duration. We demonstrate that if the particles responsible for the EMDE (but are not the dark matter in the present universe) possess self-interactions, gravothermal collapse can readily occur before the Hot Big Bang era begins.
@@ III.1 Halo formation
Once the density contrast becomes of order unity, the matter perturbations become non-linear, and eq. (7) breaks down. In this non-linear regime, the particles in overdense regions collapse to form a halo. The scale factor when halos form, $a_{\rm NL}$, can be determined as the time when the linear density contrast becomes
$\Delta_{\delta}(k,a_{\rm NL})=1.686.$ (11)
Here, the left-hand side is calculated using eq. (7). Because of the stochasticity in density perturbations, halos of a given wave number $k$ would collapse at different scale factors. The mean of the collapse scale factors is given by $a_{\rm NL}(k)$ derived from the above equation. In this study, we are primarily interested in order-of-magnitude estimates of the PBH abundance and hence we ignore the stochasticity. The dimensionless curvature power spectrum $\Delta^{2}_{\mathcal{R}}(k)$ is roughly scale-invariant and is of order $\sim 10^{-9}$ [44]. Hence we find [^1] Most results in this paper have a precision only at the order of magnitude level, as we indicate by $\sim$. However, we report one significant digit in the intermediate steps to avoid errors from being magnified by large powers in the final results.
@@
The parameters of the NFW profile can be connected to those of the background cosmology, by identifying $M_{\mathrm{halo}}$ in eq. (13) with the virial mass, and also assuming that within the virial radius (i.e. radius enclosing $M_{\mathrm{halo}}$), the mean density is $200\rho_{\phi}(a_{\rm NL})$. We refer to the virial radius upon halo formation as the halo radius, which is written in terms of the wave number of the collapsing mode as
$r_{\rm halo}=200^{-1/3}\frac{a_{\rm NL}}{k}.$ (16)
Its ratio to the scale radius defines the concentration upon halo formation,
@@
In the literature, the virial radius is defined such that it increases with time (proportional to the scale factor). Consequently, the concentration factor evolves with time even if the halo central profile is stationary [46]. However, we stress that in this paper, we use $c$ to denote the concentration at the time of the halo formation.
@@
We define the virial velocity of a halo upon its formation as
$v_{\rm vir}^{2}\equiv\frac{GM_{\rm halo}}{r_{\rm halo}},$ (20)
and also a scale velocity by $v_{s}\equiv r_{s}\sqrt{4\pi G\rho_{s}}$. These are related as,
@@
In the second relation, we used the fact that the universe is matter-dominated between the horizon entry of the mode and halo formation. This, combined with (12), yields (see also [47])
@@
To connect the halo properties with the background cosmology one needs to numerically calculate $c$ through simulations. Ref. [48] has performed N-body simulations for a matter-dominated universe with a power spectrum taking a form $P(k)=(2\pi^{2}/k^{3})\Delta_{\delta}^{2}(k)\propto k$, which corresponds to the EMDE universe under consideration (cf. eqs. (7) and (10)). According to their fig. 19, the concentration upon halo formation [^2] Fig. 19 in Ref. [48] shows concentration as a function of the halo mass normalized by the mass when it becomes non-linear. Thus, $c$ can be read off from where the normalized mass is unity. is $c\approx 30$. Since we are ignoring the stochasticity in halo properties, in the rest of the paper we fix all halos to have $c=30$. In particular, with this value for $c$, one can relate the scale density in eq. (19) with the reheating scales using eq. (12) as,
@@ III.2 Gravothermal time scale
After the halo is formed, self-scatterings cause a redistribution of the $\phi$ particles and give rise to a uniform-density core at the centre of the halo. Considering that the halo remains isolated, the redistribution occurs on a relaxation time scale determined by the mean time between the particle collisions, $\tau_{r}\sim 1/(\rho v\sigma)$ [24, 26, 32, 25, 27]. Here, $\sigma$ is the self-scattering cross section per unit mass, which for the model in eq. (5) is
@@
As the rate of gravothermal evolution becomes faster with time, most of the increase in $\rho_{c}$ occurs near $t_{\rm GC}$. Consequently, the formation of a final stable compact object (e.g. a black hole) is expected to take place also at $\sim t_{\rm GC}$.
@@
×$\times\left(\frac{a_{\mathrm{rh}}/a_{\mathrm{i}}}{10^{10}}\right)^{5/2},\ \ \frac{a_{\rm NL}}{a_{\rm i}}\bigg{]}.$ (32)
@@
[…]
@@
In Fig. 1, we illustrate when perturbation modes with fixed comoving wavelengths enter the horizon (blue), become non-linear (purple), and undergo a gravothermal collapse (red). In other words, the lines show respectively $a_{\mathrm{hor}}$, $a_{\rm NL}$, and $a_{\rm GC}$ as functions of $k^{-1}$. The EMDE parameters are chosen here such that the lines are well separated. One sees that for larger wavelengths, the separation $a_{\mathrm{GC}}/a_{\mathrm{NL}}$ between gravothermal collapse and halo formation increases. This is a consequence of the fact that halos corresponding to larger wavelengths form later and thus have smaller initial central densities, which in turn lead to slower gravothermal evolutions due to $\tau_{r}\propto 1/\rho$.
@@ III.3 Core evolution
The increase of $v_{c}$ along with the decrease of $M_{c}$ during the gravothermal evolution implies that $32<\alpha<$ 3.
@@
The transition between the LMFP and SMPF regimes takes place when $\tau_{\mathrm{dyn},c}/\tau_{r,c}=1$. Hence, using the last scaling relation in eq. (45) with $\alpha=2.190$, the ratio between the core radii at the onset of the SMFP and LMFP evolutions is found to be
@@ IV Consequences of gravothermal collapse
In this section, we discuss the final objects obtained after the gravothermal collapse. We first consider black hole formation at the end of gravothermal evolution and discuss their final masses. Next, we highlight how annihilations or pressure induced by the particles’s self-interactions can become important during gravothermal evolution and possibly impede black hole formation.
@@ IV.1.1 Efficiency of gravothermal accretion
We can roughly estimate the maximum mass a black hole can accrete from energy conservation considerations. For a virialized halo, the total energy is equivalent to the kinetic energy multiplied by $-1$. Hence, if the total energy is conserved, so is the kinetic energy. We thus equate the total kinetic energy of the particles making up the initial NFW halo and that of the halo at a later time when a black hole is present at the center,
@@
We estimate $K_{\mathrm{BH}}$ by crudely modelling a black hole as a collection of particles with a velocity that triggers the relativistic instability, i.e. $v\sim 1/3$. This yields
@@
[…]
@@
which for $c=30$ gives $F\approx 2$. Substituting eq. (56) into eq. (55) and using $v_{\rm vir}\sim 0.5\times 10^{-2}$, we find the maximum value of $\eta$ to be [^9] One can check that $\eta_{\mathrm{max}}>\eta_{\mathrm{min}}$ holds for arbitrary $c$, given that the halo is initially opcitally thin. This provides a sanity check on our computation of $\eta_{\rm max}$. We should also remark that the computation breaks down for an extremely large $c$ such that the expression for $\eta_{\mathrm{max}}$ exceeds unity. Since there, the mean-square velocity $\langle v^{2}\rangle=F(c)v_{\mathrm{vir}}^{2}$ (cf. (56)) is also large, relativistic corrections cannot be ignored.
@@ IV.1.2 Mass limits
×$\times\left(\frac{a_{\mathrm{rh}}/a_{\mathrm{i}}}{10^{10}}\right)^{-1.1}.$ (60)
We remind the reader that we have been setting the concentration of halos upon their formation to the mean value $c=30$. If one instead keeps $c$ arbitrary, then the above upper limits on $M_{\mathrm{BH}}$ for $\eta=\eta_{\mathrm{min}}$ and $\eta_{\mathrm{max}}$ scale respectively as $\propto c^{2.9}$ and $c^{1.8}$, up to logarithmic corrections. [^11] Here we also ignore the $c$-dependence of $\eta_{\mathrm{max}}$, since it is weak for $c\lesssim 10^{2}$. Thus, rarer halos with larger $c$ values can produce significantly heavier black holes. For brevity, in this study, we assume all halos have average properties and defer the study of variance in halos for the future.
@@
×$\quad\times\left(\frac{a_{\mathrm{rh}}/a_{\mathrm{i}}}{10^{10}}\right)^{-1.5}.$ (62)
@@ IV.2 Cannibal stars
So far, we have considered that the gravothermal collapse ultimately leads to the formation of a black hole by supposing the absence of any internal heat sources. However, 4-point interactions, which give rise to $2\to 2$ scatterings, generically also allow for $4\rightarrow 2$ annihilation reactions inside the nonrelativictic core. (The inverse process is energetically forbidden.) These cannibal reactions convert the rest-mass energy into kinetic energy and can provide support against gravitational collapse, similar to how nuclear fusion supports a star. Below we discuss the conditions for the formation of such “cannibal stars,” and also hypothesize on their possible collapse into black holes after accretion.
@@ V PBH abundance
While the halos as well as the various stars arising from their gravothermal collapse disappear as the $\phi$ particles decay, black holes can survive after the end of EMDE. In this section, we estimate the relic PBH abundance. Since we have been discussing the evolution of isolated halos, we focus on halos that do not merge into a bigger halo before the gravothermal collapse.
@@ V.1 PBHs relative to dark matter
In the second line of eq. (99), we highlighted the terms that can depend on the wave number. By writing $k$ in terms of $M_{\mathrm{BH}}$ using the results in Section IV.1.1, we can express the PBH fraction as an explicit function of the PBH mass. For this purpose, let us evaluate $a_{\mathrm{GC}}$ using the first term in the square brackets in eq. (III.2) by considering PBHs formed in halos that were initially collisionless (cf. discussions around eq. (35)). Moreover, we calculate the redshift at reheating by assuming entropy conservation from reheating onwards as,
@@
×$\times\left(\frac{M_{\mathrm{BH}}}{10^{16}~{}\mathrm{g}}\right)^{-2.5}.$
@@
$f_{\mathrm{BH}}[\eta_{\rm max}]\sim{}$ $10^{6}~{}\lambda^{1.3}\left(\frac{\eta_{\rm max}}{10^{-3}}\right)^{1.7}\left(\frac{a_{\mathrm{rh}}/a_{\mathrm{i}}}{10^{10}}\right)^{-1.5}$ (102)
×$\times\left(\frac{T_{\mathrm{rh}}}{100~{}\mathrm{MeV}}\right)^{-1}\left(\frac{M_{\mathrm{BH}}}{10^{16}~{}\mathrm{g}}\right)^{-0.67}.$
@@ V.2 Evaporating PBHs relative to radiation
$\kappa_{\rm BH}[\eta_{\rm min}]\sim$ $10^{-16}\,\lambda^{5.0}\left(\frac{a_{\mathrm{rh}}/a_{\mathrm{i}}}{10^{15}}\right)^{-5.6}$ (106)
×$\times\left(\frac{T_{\mathrm{rh}}}{100~{}\mathrm{MeV}}\right)^{-6.4}\left(\frac{M_{\mathrm{BH}}}{10^{9}~{}\mathrm{g}}\right)^{-0.98},$
$\kappa_{\rm BH}[\eta_{\rm max}]\sim$ $10^{-3}\,\lambda^{1.3}\left(\frac{\eta_{\rm max}}{10^{-3}}\right)^{1.7}\left(\frac{a_{\mathrm{rh}}/a_{\mathrm{i}}}{10^{15}}\right)^{-1.5}$ (107)
×$\times\left(\frac{T_{\mathrm{rh}}}{100~{}\mathrm{MeV}}\right)^{-1}\left(\frac{M_{\mathrm{BH}}}{10^{9}~{}\mathrm{g}}\right)^{0.83}.$

### v3

@@ Abstract
Very little is known about the cosmological history from after the end of inflation until Big Bang Nucleosynthesis. Various well-motivated models predict that the universe could have undergone a period of matter domination in this early epoch. We demonstrate that if the particles causing matter domination have self-interactions, they can form halos that undergo a gravothermal collapse. We thus propose a novel scenario for the formation of primordial black holes, which in particular can lie within the asteroid-mass range. We also find that it is not only black holes that can form in the aftermath of a gravothermal evolution. We show that number-changing annihilations of the particles can create sufficient heat to halt the gravothermal evolution, thus forming a “cannibal star”. Likewise, the pressure from the particle’s repulsive self-interactions can form a boson star during a gravothermal evolution. These stars can also eventually collapse into black holes. Thus, our study highlights that structure formation in the early universe can have a rich phenomenology.
@@ I Introduction
The period of matter domination in the standard Hot Big Bang history lasts about eight $e$-folds. An EMDE, on the other hand, can have a much longer duration. We demonstrate that if the particles responsible for the EMDE (but are not the dark matter in the present universe) possess self-interactions, gravothermal collapse can readily occur before the Hot Big Bang era begins.
@@ III.1 Halo formation
Once the density contrast becomes of order unity, the matter perturbations become non-linear, and eq. (7) breaks down. In this non-linear regime, the particles in overdense regions collapse to form a halo. We define $a_{\mathrm{NL}}$ as the scale factor when the linear density contrast becomes
$\Delta_{\delta}(k,a_{\rm NL})=1.686,$ (11)
with the left-hand side calculated using eq. (7). This can be considered as the moment when halos corresponding to the wave number $k$ form. [^1] More precisely, $a_{\mathrm{NL}}(k)$ defined via eq. (11) corresponds to the time when the number of halos with mass $M_{\mathrm{halo}}(k)$ (cf. eq. (13)) is around its maximum. See Appendix B. Here we remark that the stochasticity in density perturbations induces halos of a given $k$ to collapse at different times, and $a_{\rm NL}(k)$ defined above denotes the mean of the collapse scale factors. In this study, we are primarily interested in order-of-magnitude estimates of the PBH abundance and hence we ignore the stochasticity. The dimensionless curvature power spectrum $\Delta^{2}_{\mathcal{R}}(k)$ is roughly scale-invariant and is of order $\sim 10^{-9}$ [44]. Hence we find [^2] Most results in this paper have a precision only at the order of magnitude level, as we indicate by $\sim$. However, we report one significant digit in the intermediate steps to avoid errors from being magnified by large powers in the final results.
@@
The parameters of the NFW profile can be connected to those of the background cosmology, by identifying $M_{\mathrm{halo}}$ in eq. (13) with the virial mass, and also assuming that within the virial radius (i.e. radius enclosing $M_{\mathrm{halo}}$), the mean density is $200\rho_{\phi}(a_{\rm NL})$. We define the halo radius as the virial radius of the halo of mass $M_{\rm halo}(k)$ at $a_{\rm NL}(k)$, i.e.
$r_{\rm halo}(k)=200^{-1/3}\frac{a_{\rm NL}(k)}{k}.$ (16)
Its ratio to the scale radius defines the concentration at $a_{\rm NL}$,
@@
In the literature, the virial radius is defined such that it increases with time (proportional to the scale factor). Consequently, the concentration factor evolves with time even if the halo central profile is stationary [46]. However, we stress that in this paper, we use $c$ to denote the concentration at $a_{\rm NL}$.
@@
We define the virial velocity of a halo as
$v_{\rm vir}^{2}\equiv\frac{GM_{\rm halo}}{r_{\rm halo}}.$ (20)
We also define the scale velocity as $v_{s}\equiv r_{s}\sqrt{4\pi G\rho_{s}}$. These are related as,
@@
In the second relation, we used the fact that the universe is matter dominated between the horizon entry of the mode and halo formation. This, combined with (12), yields (see also [47])
@@
To connect the halo properties with the background cosmology one needs to numerically calculate $c$ through simulations. Ref. [48] has performed N-body simulations for a matter-dominated universe with a power spectrum taking a form $P(k)=(2\pi^{2}/k^{3})\Delta_{\delta}^{2}(k)\propto k$, which corresponds to the EMDE universe under consideration (cf. eqs. (7) and (10)). According to their fig. 19, the concentration of halos at their formation is $c\approx 30$. [^3] Fig. 19 in Ref. [48] shows concentration as a function of the halo mass normalized by the mass when it becomes non-linear. Thus, $c$ can be read off from where the normalized mass is unity. Since we are ignoring the stochasticity in halo properties, in the rest of the paper we fix all halos to have $c=30$. In particular, with this value for $c$, one can relate the scale density at halo formation in eq. (19) with the reheating scales using eq. (12) as,
@@ III.2 Gravothermal time scale
After the halo is formed, self-scatterings cause a redistribution of the $\phi$ particles and give rise to a uniform-density core at the center of the halo. Considering that the influences of other halos (such as their tidal force) are negligible, the redistribution occurs on a relaxation time scale determined by the mean time between the particle collisions, $\tau_{r}\sim 1/(\rho v\sigma)$ [24, 26, 32, 25, 27]. Here, $\sigma$ is the self-scattering cross section per unit mass, which for the model in eq. (5) is
@@
As the rate of gravothermal evolution becomes faster with time, most of the increase in $\rho_{c}$ occurs near $t_{\rm GC}$. Consequently, the formation of a final stable compact object (e.g. a black hole) is expected to take place also at $\sim t_{\rm GC}$. [^5] We stress that this discussion assumes the halo to remain isolated from external tidal forces until it collapses. Since most halos undergo continual mergers and accretion, it is only rare halos that can stay isolated for significant periods; we estimate the abundance of such halos in Section V.
@@
$\times\left(\frac{a_{\mathrm{rh}}/a_{\mathrm{i}}}{10^{10}}\right)^{5/2},\ \ \frac{a_{\rm NL}}{a_{\rm i}}\bigg{]}.$ (32)
@@
$\times\left(\frac{a_{\mathrm{rh}}/a_{\mathrm{i}}}{10^{10}}\right)^{3/8},\ \ 4\times 10^{2}\bigg{]}.$ (34)
@@
In Fig. 1, we illustrate when perturbation modes with fixed comoving wavelengths enter the horizon (blue), become non-linear (purple), and undergo a gravothermal collapse (red). In other words, the lines show respectively $a_{\mathrm{hor}}$, $a_{\rm NL}$, and $a_{\rm GC}$ as functions of $k^{-1}$. The EMDE parameters are chosen here such that the lines are well separated. One sees that for larger wavelengths, the separation $a_{\mathrm{GC}}/a_{\mathrm{NL}}$ between gravothermal collapse and halo formation increases. This is a consequence of the fact that longer-wavelength modes form halos later and with lower central densities, which in turn lead to slower gravothermal evolutions due to $\tau_{r}\propto 1/\rho$.
@@ III.3 Core evolution
The increase of $v_{c}$ along with the decrease of $M_{c}$ during the gravothermal evolution implies that $2<\alpha<3$.
@@
The transition between the LMFP and SMFP regimes takes place when $\tau_{\mathrm{dyn},c}/\tau_{r,c}=1$. Hence, using the last scaling relation in eq. (45) with $\alpha=2.190$, the ratio between the core radii at the onset of the SMFP and LMFP evolutions is found to be
@@ IV Consequences of gravothermal collapse
In this section, we discuss the final objects obtained after the gravothermal collapse. We first consider black hole formation at the end of gravothermal evolution and discuss their final masses. Next, we highlight how annihilations or pressure induced by the particle’s self-interactions can become important during gravothermal evolution and possibly impede black hole formation.
@@ IV.1.1 Efficiency of gravothermal accretion
We can roughly estimate the maximum mass a black hole can accrete from energy conservation considerations. For a virialized halo, the total energy is equivalent to the kinetic energy multiplied by 1-1. Hence, if the total energy is conserved, so is the kinetic energy. We thus equate the total kinetic energy of the particles making up the initial NFW halo and that of the halo at a later time when a black hole is present at the center,
@@
We estimate $K_{\mathrm{BH}}$ by crudely modeling a black hole as a collection of particles with a velocity that triggers the relativistic instability, i.e. $v\sim 1/3$. This yields
@@
$\begin{split}K_{\rm NFW}=-\frac{1}{2}U(r_{\mathrm{halo}})&=2\pi G\int^{r_{\mathrm{halo}}}_{0}M(r)\rho(r)r\,dr\\ &=\frac{1}{2}F(c)M_{\mathrm{halo}}v_{\rm vir}^{2}.\end{split}$ (56)
@@
which for $c=30$ gives $F\approx 2$. Substituting eq. (56) into eq. (55) and using $v_{\rm vir}\sim 0.5\times 10^{-2}$, we find the maximum value of $\eta$ to be [^11] One can check that $\eta_{\mathrm{max}}>\eta_{\mathrm{min}}$ holds for arbitrary $c$, given that the halo is initially optically thin. This provides a sanity check on our computation of $\eta_{\rm max}$. We should also remark that the computation breaks down for an extremely large $c$ such that the expression for $\eta_{\mathrm{max}}$ exceeds unity. Since there, the mean-square velocity $\langle v^{2}\rangle=F(c)v_{\mathrm{vir}}^{2}$ (cf. (56)) is also large, relativistic corrections cannot be ignored.
@@ IV.1.2 Mass limits
$\times\left(\frac{a_{\mathrm{rh}}/a_{\mathrm{i}}}{10^{10}}\right)^{-1.1}.$ (60)
We remind the reader that we have been setting the concentration of halos upon their formation at $a_{\rm NL}$ to the mean value $c=30$. If one instead keeps $c$ arbitrary, then the above upper limits on $M_{\mathrm{BH}}$ for $\eta=\eta_{\mathrm{min}}$ and $\eta_{\mathrm{max}}$ scale respectively as $\propto c^{2.9}$ and $c^{1.8}$, up to logarithmic corrections. [^13] Here we also ignore the $c$-dependence of $\eta_{\mathrm{max}}$, since it is weak for $c\lesssim 10^{2}$. Thus, rarer halos with larger $c$ values can produce significantly heavier black holes. For brevity, in this study, we assume all halos have average properties and defer the study of variance in halos for the future.
@@
$\quad\times\left(\frac{a_{\mathrm{rh}}/a_{\mathrm{i}}}{10^{10}}\right)^{-1.5}.$ (62)
@@ IV.2 Cannibal stars
So far, we have considered that the gravothermal collapse ultimately leads to the formation of a black hole by supposing the absence of any internal heat sources. However, 4-point interactions, which give rise to $2\to 2$ scatterings, generically also allow for $4\rightarrow 2$ annihilation reactions inside the nonrelativistic core. (The inverse process is energetically forbidden.) These cannibal reactions convert the rest-mass energy into kinetic energy and can provide support against gravitational collapse, similar to how nuclear fusion supports a star. Below we discuss the conditions for the formation of such “cannibal stars,” and also hypothesize on their possible collapse into black holes after accretion.
@@ V PBH abundance
While the halos as well as the various stars arising from their gravothermal collapse disappear as the $\phi$ particles decay, black holes can survive after the end of EMDE. In this section, we estimate the relic PBH abundance. Since we have been discussing the evolution of isolated halos, i.e. halos that are not significantly disrupted by external tidal forces, here we focus on the rare halos that do not merge into a bigger halo or undergo significant accretion before the gravothermal collapse.
@@
In the above abundance estimate, we are implicitly assuming that the halos with mass $M_{\rm halo}(k)$ at $a_{\mathrm{GC}}(k)$ were formed at $a_{\mathrm{NL}}(k)$, and hence have stayed isolated for a collapse time. In principle halos with mass $M_{\mathrm{halo}}(k)$ can newly form even long after $a_{\mathrm{NL}}(k)$, however, we expect the number of such halos to be negligibly tiny (cf. footnote 27).
@@ V.1 PBHs relative to dark matter
In the second line of eq. (99), we highlighted the terms that can depend on the wave number. By writing $k$ in terms of $M_{\mathrm{BH}}$ using the results in Section IV.1.1, we can express the PBH fraction as an explicit function of the PBH mass. For this purpose, let us evaluate $a_{\mathrm{GC}}$ using the first term in the square brackets in eq. (III.2) by considering PBHs formed in halos that were initially collisionless (cf. discussions around eq. (35)). Moreover, we calculate the redshift at reheating by assuming entropy conservation from reheating onward as,
@@
$\times\left(\frac{M_{\mathrm{BH}}}{10^{16}~\mathrm{g}}\right)^{-2.5}.$
@@
[…]
$\times\left(\frac{T_{\mathrm{rh}}}{100~\mathrm{MeV}}\right)^{-1}\left(\frac{M_{\mathrm{BH}}}{10^{16}~\mathrm{g}}\right)^{-0.67}.$
@@ V.2 Evaporating PBHs relative to radiation
[…]
$\times\left(\frac{T_{\mathrm{rh}}}{100~\mathrm{MeV}}\right)^{-6.4}\left(\frac{M_{\mathrm{BH}}}{10^{9}~\mathrm{g}}\right)^{-0.98},$
[…]
$\times\left(\frac{T_{\mathrm{rh}}}{100~\mathrm{MeV}}\right)^{-1}\left(\frac{M_{\mathrm{BH}}}{10^{9}~\mathrm{g}}\right)^{0.83}.$